Gini Coefficient.
A = 0.2
Determine the Gini Coefficient...
1
Expert's answer
2016-05-08T12:35:03-0400
A = 0.2, B = 0.28, where A is the area between the diagonal and the Lorenz curve, B is the rest area under the diagonal.
Gini Coefficient G = A/(A + B) = 0.2/(0.2 + 0.28) = 0.476, so as G is lower than 0.5 and there an inequality.